factor and solve equation
(x+5)(x-3)=0

Answer by stanbon(75887)

factor and solve equation
(x+5)(x-3)=0
---
Since the product is zero, one of those factors must be 0.
---
x+5 = 0 or x-3 = 0
x = -5 or x = 3
=========================
cheers,
Stan H.